“BLACKFRIDAY” company is planning an expansion of its existing production capacity. The firm hired you as a consultant for the expansion project. Since you are a savvy project manager, you first decided to estimate the firm’s cost of capital based on the available data.

Data:

  • Tax Rate: 40%
  • Bond: Coupon rate 12%, Maturity Years 15, Present value $1150
  • Preferred Stock: Dividend rate 10%, Par Value $100, Present Value $111 Common Stock: Market price $50, D0=$4.20, Dividend growth 5%, Beta 1.2, Treasury Bond yield 7%, Market risk premium 6%. When the firm uses Bond-yield+Premium method, the risk premium is 4%.
  • Capital structure of “BLACKFRIDAY” is as follows;
    • Debt 30%, Common Equity 60%, Preferred Stock 10%
  1. If the flotation cost of new stock issue is 10%, what is the estimated cost of equity considering the 10% flotation cost under DCF method you calculated in question 4 above?
  2. What are the components of the WACC you calculated above to be blamed for higher WACC? What can be done to reduce the WACC further?
  3. Should you use this WACC for all the projects? Why and why not?
Answer #1
1] Before tax cost of debt = YTM
YTM using an online YTM calculator = 10.03%
After tax cost of debt = 10.03%*(1-40%) = 6.02%
2] Cost of preferred stock = 10/111 = 9.01%
3] Cost of equity:
a] Per constant dividend growth formula = 4.20*1.05/50+0.05 = 13.82%
b] Per CAPM = 7%+1.2*6% = 14.20%
c] Per bond yield+premium method = 10.03%+4.00% = 14.03%
d] Average cost of equity = (13.82%+14.20%+14.03%)/3 = 14.02%
4] WACC = 6.02%*30%+9.01%*10%+14.02%*60% = 11.12%
5] Cost of new equity under DCF = 4.20*1.05/(50-50*10%)+0.05 = 14.80%
6] The component that makes WACC higher is equity, as its cost
is high and also its proportion in the total capital.
7] WACC can be reduced by increasing the % of debt and % of
preference capital to the maximum extent possible.
8] WACC should be used only for those projects that have risks
similar to that of the existing business of the firm. Otherwise
the firm will end up accepting projects that ought to have
been rejected or rejecting projects that ought to have been
accepted.